Is the term (x+2) a factor of the polynomial shown below? f(x)=x^4+5x^3+10x^2+20x+24


1
Expert's answer
2021-11-17T17:22:40-0500
"f(x)=x^4+5x^3+10x^2+20x+24"


"f(-2)=(-2)^4+5(-2)^3+10(-2)^2+20(-2)+24"

"=16-40+40-40+24=0"

Then the term "(x+2)" is a factor of the polynomial.


"f(x)=x^4+5x^3+10x^2+20x+24"

"f(x)=x^4+2x^3+3x^3+6x^2+4x^2+8x+12x+24"

"=x^3(x+2)+3x^2(x+2)+4x(x+2)+12(x+2)"

"=(x+2)(x^3+3x^2+4x+12)"

Therefore the term "(x+2)" is a factor of the polynomial.